OpenStudy (anonymous):

What is the solution of the equation? -10 +√x+8=-4

OpenStudy (anonymous):

A. -2 B. 28 C. 36 D. 44

OpenStudy (anonymous):

@Embryo @mathslover @lolo5656 @HeroOfTimeAndSpace

OpenStudy (anonymous):

Okay, first you have to get x by itself.

OpenStudy (anonymous):

is the equation: \[-10 + \sqrt{x+8}=-4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay so isolate x. First add ten to both sides. \[\sqrt{x+8}= 6\]

OpenStudy (anonymous):

Then, you have to square everything.

OpenStudy (anonymous):

i tried it on mathway.com and it says no solution. ugh i just hate math, i really need to finish these exams and it feels impossible.

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

sorry was afk

OpenStudy (anonymous):

after squaring the equation it should be \[x + 8 = \]

OpenStudy (anonymous):

x + 8 = 36

OpenStudy (anonymous):

Now just isolate x. Subtract 8 from both sides. :)

OpenStudy (anonymous):

so its B?

OpenStudy (anonymous):

x = 28. do you get how I did that?

OpenStudy (anonymous):

yes! thank you

OpenStudy (anonymous):

Yeah it's B, but you get how to do it right?

OpenStudy (anonymous):

yes! thank you <3 @jcpd910

OpenStudy (anonymous):

Np, if you have more questions just @jcpd910 me, but don't post more than one question on the same thread, like a lot of people do. :D

OpenStudy (anonymous):

could you help with one more @jcpd910

OpenStudy (anonymous):

Post it as a new question and I will. :)